OpenStudy (anonymous):

Identify the solution(s) of the system of equations, if any. 3x+2y=4 5x-4y=3

OpenStudy (anonymous):

eq 1 *-2 and add it to eq 2

OpenStudy (anonymous):

no eq 1 * 2 sorry

OpenStudy (the_fizicx99):

You can use the substitution method or the elimination method. Which do you prefer?

OpenStudy (anonymous):

5x + 6x=3+8 x=1

OpenStudy (anonymous):

I tend to do the elimination method more

OpenStudy (anonymous):

y = .5

OpenStudy (anonymous):

(1, .5)

OpenStudy (the_fizicx99):

Ok,  Now add that, the x;s cancel out